OpenStudy (anonymous):

Evan is painting a stool that has a cylindrical seat and four rectangular prism legs. The seat has a diameter of 8 inches and a height of 2 inches. Each leg is 2 inches by 2 inches by 10 inches. How much area will Evan have to paint? Use 3.14 for pi . Show your work.

OpenStudy (pfenn1):

The area of a cylinder is given by\[A_{cylinder}=2\pi r^2+2\pi rh\]where r is the radius and h is the height. The area of each rectangular prism (leg) is given by \[A_{leg}=2(wh+lw+hl)\]where h is the height of the leg, l is the length and w is the width. In this case, the width and length are both 2 inches so in this case\[A_{leg}=2(2w^2+wh)\]